OpenStudy (anonymous):

expand each logarithm: 1. log[base 8] 8 ( square root 3a^5) 2.log (s(square root 7) / t^2) 3.log [base b] 1/x Please help !

OpenStudy (anonymous):

I think if you have a look at the log rules you will be able to solve easily. Am I correct in saying that.

OpenStudy (anonymous):

yea i looked at them , bt i jus want to see if my answers are right ... like for the fist one i got: 2 log a + 3 log b - 4 log c

OpenStudy (anonymous):

is that correct ?

OpenStudy (anonymous):

yes that would be correct

OpenStudy (anonymous):

:) okay bt i am now i am not sure about the second one ...i got : 1/2 log 2 + 1/2 log x - 1/2 log y

OpenStudy (anonymous):

s that answer right ?

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

is the square root for both the x and y term or just the x

OpenStudy (anonymous):

both the x and the y

OpenStudy (anonymous):

bt it is 2x over y and the square root is for all

OpenStudy (anonymous):

you answer is also correct.

OpenStudy (anonymous):

so does that mean \[\sqrt{?} = 1/2\]

OpenStudy (anonymous):

yes, the square root becomes a half

OpenStudy (anonymous):

oh okay thank you :) bt how about number 3

OpenStudy (anonymous):

\[\log_{8} 8 \sqrt{3a ^{5}}\]

OpenStudy (anonymous):

idk how to solve it

OpenStudy (anonymous):

 sorry I have difficulty in writing subscripts